# Projector, Neural, and Tensor-Network Representations of

### 08 Apr 2026 v1

@@ I Introduction
Representing many-body quantum states efficiently has been a central theme of condensed matter theory for nearly a century. Starting from the Hartree–Fock approximation, successive developments such as Jastrow-correlated wavefunctions, quantum and variational Monte Carlo, exact diagonalization, and density-matrix renormalization group (DMRG) have all contributed essential insights into strongly correlated states of fermions and bosons. With the advent of modern AI and transformer-based methods, the demand for compact yet accurate descriptions of correlated quantum states has only intensified, together with new opportunities to construct such descriptions [6, 35]. In particular, representing quantum many-body wavefunctions in the spirit of neural networks—most notably via restricted Boltzmann machines (RBM)—has led to the notion of neural quantum states (NQS) [6, 13, 12, 14, 15, 11, 8, 18, 33, 17, 7, 36, 21, 9, 25] which has proven capable of exactly encoding paradigmatic states such as the one-dimensional cluster state and the two-dimensional toric code [13, 12, 14, 8, 9]. In this work, we revisit and systematically generalize the NQS formulation of one-dimensional ${\mathbb{Z}}_{N}$ cluster states, focusing on symmetry-protected topological (SPT) order [10, 29, 26] and its multipolar extensions [16, 20, 19, 34, 24, 28, 4, 23, 2], and clarify its relation to matrix product state (MPS) and more general tensor product state (TPS) constructions. In the process, we develop another efficient way to represent various cluster states based on a local projector, which we name the $P$-representation.
@@
Much progress has been made since the first NQS representation of the ${\mathbb{Z}}_{2}$ cluster state appeared [13, 12, 14, 8]. The ${\mathbb{Z}}_{2}$ cluster state, originally introduced as a universal resource for measurement-based quantum computation [27], has been generalized to ${\mathbb{Z}}_{N}$ cluster states with arbitrary integer $N>2$ and to its variants protected by modulated symmetries—such as dipolar, quadrupolar, and exponential symmetries—which define a broader class of modulated SPT (mSPT) states [16, 20, 28, 22, 37, 4, 2] along with the identification of non-invertible Kramers-Wannier symmetries possessed by the various cluster models [30, 31, 5, 19]. Motivated by these developments, we return to the cluster-state NQS framework with two main goals. First, we extend the NQS construction to uniform ${\mathbb{Z}}_{N}$ and dipole-symmetric ${\mathbb{Z}}_{N}$ cluster models. This is made possible by providing the first closed-form RBM weights for general ${\mathbb{Z}}_{N}$ cluster states. Second, we sharpen the understanding of symmetry fractionalization and and other symmetry aspects of these SPT wavefunctions. Concretely, we construct explicit NQS representations for the ${\mathbb{Z}}_{N}$ cluster state and for two types of dipole-conserving ${\mathbb{Z}}_{N}$ cluster states (type-I and type-II), derive closed-form interaction weight matrices for ${\mathbb{Z}}_{N}$ variables, and show how different ways of splitting the weight matrices yield either an MPS or, in the dipolar case, a TPS with three virtual indices per site. In accomplishing these goals, we introduce the $P$-representation and make extensive use of it as a unifying bridge between NQS and tensor-network-type representations. Using the $P$-reprensentation also helps streamlining proofs of push-through conditions, symmetry fractionalization, and writing down matrix-product-operator (MPO) representation of the ${\mathbb{Z}}_{N}$ Kramers–Wannier (KW) operator. Finally we show both analytically and numerically by performing DMRG calculations that certain dipolar SPT states naturally require TPS as the more efficient way to represent its many-body state. As a by-product of our investigation, we propose the interpretation of the KW operation as the dipolar version of the discrete Fourier transformation and give an intuitive explanation for the non-invertibility of the KW operation.
@@ II.1 Definition and basic properties
generating the NN-fold degeneracy at each edge. The corresponding eigenstates can be constructed explicitly as, e.g. $|\Psi(s_{1},s_{L})\rangle=L_{1}^{s_{1}}R_{2}^{s_{L}}|\Psi(0,0)\rangle$, using $\{s_{1},s_{L}\}$ as quantum numbers labeling each ground state.
@@ II.2 CS wavefunction in $P$, NQS, and MPS representations
For even NN, it suffices to choose (a,b,c)=(1/4,1/2,1)(a,b,c)=(-1/4,-1/2,-1). For derivation, see App. B. We can write (II.2) as a matrix relation:
@@ III.1 Type-I dipolar cluster state
where $L$ must be divisible by NN for a closed chain to unambiguously define the dipole symmetry operator $D$. Its unique ground state on a periodic chain is [16]
@@ III.4 Representational aspects of NQS and TPS
To compare the TPS with the conventional MPS, we employ DMRG to numerically obtain the MPS ground state of the dCSII Hamiltonian. The resulting bond dimension of the MPS tensor $A^{s}$ scales as $\chi=N^{2}$. Meanwhile, the TPS tensors in Eq. (III.2) each carry three virtual indices of dimension NN, so that the local tensor has $N^{4}$ elements — a factor of NN smaller than the $N^{5}$ elements of the MPS tensor with bond dimension $N^{2}$. Thus, at the level of local tensors, the TPS provides a more compact representation of the dCSII ground state.
We note, however, that the compactness of local tensors does not directly translate into computational efficiency for the full network contraction. In the MPS representation, the cost of computing the norm or local observables scales as $O(N^{7}L)$. In the TPS representation, the third virtual index $\gamma_{j}$ connects non-nearest-neighbor sites and introduces loops into the tensor network. Even with an optimized contraction ordering, these loops incur an additional factor of NN, leading to a contraction cost of $O(N^{8}L)$. Thus, despite the smaller local tensor size, exact contraction of the TPS is a factor of NN more expensive than MPS. The advantage of the NQS/TPS construction for the dCSII state is therefore primarily representational — it reveals the natural tensor structure dictated by the modulated symmetries — rather than computational.
@@ IV Kramers-Wannier operation as dipolar Fourier transform and its MPO
Now the target space has two constraints, in the form of the total “charge” and the total “dipole” conservation: $\sum_{j}q_{j}=\sum_{j}jq_{j}=0$  [^1] More precisely, the total dipole moment must be zero modulus NN, assuming that the length of the chain $L$ is divisible by NN.. The transformed function is invariant under both $s_{j}\rightarrow s_{j}+1$ and $s_{j}\rightarrow s_{j}+j$, i.e. resides in the space where both $C$ and $D=\prod_{j}(X_{j})^{j}$ are effectively one. The quadrupole Fourier transformation appears as the KW operator implementing the NIS of the dCSII - see App. F.
@@ V Summary and discussion
Perturbations of the cluster Hamiltonians will in general deform the corresponding ground states. In the present language, such deformations can be viewed as renormalizations of the interaction matrix $\bm{\Omega}$, or the weight matrix $\bf W$}, in analogy to how a Slater determinant is modified by a Jastrow factor or how fixed-point tensors are renormalized away from exact SPT fixed points. It is plausible that these renormalizations can be implemented efficiently using modern AI-based techniques for learning many-body wavefunctions [38], or that the nonlinearity structure of the weight functions is tied to the entanglement content of the state [25]. Exploring these directions—both for perturbed cluster models and for more general modulated SPT states—remains an interesting avenue for future research. Representing a two-dimensional topological state such as the toric code and SPT states with ${\mathbb{Z}}_{N}$ physical degrees of freedom is another exciting avenue of future research.

### 14 Apr 2026 v2

@@ I Introduction
Representing many-body quantum states efficiently has been a central theme of condensed matter theory for nearly a century. Starting from the Hartree–Fock approximation, successive developments such as Jastrow-correlated wavefunctions, quantum and variational Monte Carlo, exact diagonalization, and density-matrix renormalization group (DMRG) have all contributed essential insights into strongly correlated states of fermions and bosons. With the advent of modern AI and transformer-based methods, the demand for compact yet accurate descriptions of correlated quantum states has only intensified, together with new opportunities to construct such descriptions [6, 39]. In particular, representing quantum many-body wavefunctions in the spirit of neural networks—most notably via restricted Boltzmann machines (RBM)—has led to the notion of neural quantum states (NQS) [6, 13, 12, 14, 15, 11, 8, 19, 18, 37, 17, 7, 40, 24, 9, 29] which has proven capable of exactly encoding paradigmatic states such as the one-dimensional cluster state and the two-dimensional toric code [13, 12, 14, 8, 9]. In this work, we revisit and systematically generalize the NQS formulation of one-dimensional ${\mathbb{Z}}_{N}$ cluster states, focusing on symmetry-protected topological (SPT) order [10, 33, 30] and its multipolar extensions [16, 23, 22, 38, 28, 32, 4, 27, 2, 26], and clarify its relation to matrix product state (MPS) and more general tensor product state (TPS) constructions. In the process, we develop another efficient way to represent various cluster states based on a local projector, which we name the $P$-representation.
@@
Much progress has been made since the first NQS representation of the ${\mathbb{Z}}_{2}$ cluster state appeared [13, 12, 14, 8]. The ${\mathbb{Z}}_{2}$ cluster state, originally introduced as a universal resource for measurement-based quantum computation [31], has been generalized to ${\mathbb{Z}}_{N}$ cluster states with arbitrary integer $N>2$ and to its variants protected by modulated symmetries—such as dipolar, quadrupolar, and exponential symmetries—which define a broader class of modulated SPT (mSPT) states [16, 23, 32, 25, 41, 4, 2, 26] along with the identification of non-invertible Kramers-Wannier symmetries possessed by the various cluster models [34, 35, 5, 22]. Motivated by these developments, we return to the cluster-state NQS framework with two main goals. First, we extend the NQS construction to uniform ${\mathbb{Z}}_{N}$ and dipole-symmetric ${\mathbb{Z}}_{N}$ cluster models. This is made possible by providing the first closed-form RBM weights for general ${\mathbb{Z}}_{N}$ cluster states. Second, we sharpen the understanding of symmetry fractionalization and and other symmetry aspects of these SPT wavefunctions. Concretely, we construct explicit NQS representations for the ${\mathbb{Z}}_{N}$ cluster state and for two types of dipole-conserving ${\mathbb{Z}}_{N}$ cluster states (type-I and type-II), derive closed-form interaction weight matrices for ${\mathbb{Z}}_{N}$ variables, and show how different ways of splitting the weight matrices yield either an MPS or, in the dipolar case, a TPS with three virtual indices per site. In accomplishing these goals, we introduce the $P$-representation and make extensive use of it as a unifying bridge between NQS and tensor-network-type representations. Using the $P$-reprensentation also helps streamlining proofs of push-through conditions, symmetry fractionalization, and writing down matrix-product-operator (MPO) representation of the ${\mathbb{Z}}_{N}$ Kramers–Wannier (KW) operator. Finally we show both analytically and numerically by performing DMRG calculations that certain dipolar SPT states naturally require TPS as the more efficient way to represent its many-body state. As a by-product of our investigation, we propose the interpretation of the KW operation as the dipolar version of the discrete Fourier transformation and give an intuitive explanation for the non-invertibility of the KW operation.
@@ II.1 Definition and basic properties
generating the $N$-fold degeneracy at each edge. The corresponding eigenstates can be constructed explicitly as, e.g. $|\Psi(s_{1},s_{L})\rangle=L_{1}^{s_{1}}R_{2}^{s_{L}}|\Psi(0,0)\rangle$, using $\{s_{1},s_{L}\}$ as quantum numbers labeling each ground state.
@@ II.2 CS wavefunction in $P$, NQS, and MPS representations
For even $N$, it suffices to choose (a,b,c)=(1/4,1/2,1)(a,b,c)=(-1/4,-1/2,-1). For derivation, see App. B. We can write (II.2) as a matrix relation:
@@ III.1 Type-I dipolar cluster state
where $L$ must be divisible by $N$ for a closed chain to unambiguously define the dipole symmetry operator $D$. Its unique ground state on a periodic chain is [16]
@@ III.4 Representational aspects of NQS and TPS
To compare the TPS with the conventional MPS, we employ DMRG to numerically obtain the MPS ground state of the dCSII Hamiltonian. The resulting bond dimension of the MPS tensor $A^{s}$ scales as $\chi=N^{2}$. Meanwhile, the TPS tensors in Eq. (III.2) each carry three virtual indices of dimension $N$, so that the local tensor has $N^{4}$ elements — a factor of $N$ smaller than the $N^{5}$ elements of the MPS tensor with bond dimension $N^{2}$. Thus, at the level of local tensors, the TPS provides a more compact representation of the dCSII ground state.
We note, however, that the compactness of local tensors does not directly translate into computational efficiency for the full network contraction. In the MPS representation, the cost of computing the norm or local observables scales as $O(N^{7}L)$. In the TPS representation, the third virtual index $\gamma_{j}$ connects non-nearest-neighbor sites and introduces loops into the tensor network. Even with an optimized contraction ordering, these loops incur an additional factor of $N$, leading to a contraction cost of $O(N^{8}L)$. Thus, despite the smaller local tensor size, exact contraction of the TPS is a factor of $N$ more expensive than MPS. The advantage of the NQS/TPS construction for the dCSII state is therefore primarily representational — it reveals the natural tensor structure dictated by the modulated symmetries — rather than computational.
@@ IV Kramers-Wannier operation as dipolar Fourier transform and its MPO
Now the target space has two constraints, in the form of the total “charge” and the total “dipole” conservation: $\sum_{j}q_{j}=\sum_{j}jq_{j}=0$  [^1] More precisely, the total dipole moment must be zero modulus $N$, assuming that the length of the chain $L$ is divisible by $N$.. The transformed function is invariant under both $s_{j}\rightarrow s_{j}+1$ and $s_{j}\rightarrow s_{j}+j$, i.e. resides in the space where both $C$ and $D=\prod_{j}(X_{j})^{j}$ are effectively one. The quadrupole Fourier transformation appears as the KW operator implementing the NIS of the dCSII - see App. F.
@@ V Summary and discussion
Perturbations of the cluster Hamiltonians will in general deform the corresponding ground states. In the present language, such deformations can be viewed as renormalizations of the interaction matrix $\bm{\Omega}$, or the weight matrix $\bf W$}, in analogy to how a Slater determinant is modified by a Jastrow factor or how fixed-point tensors are renormalized away from exact SPT fixed points [20]. It is plausible that these renormalizations can be implemented efficiently using modern AI-based techniques for learning many-body wavefunctions [42], or that the nonlinearity structure of the weight functions is tied to the entanglement content of the state [29]. Exploring these directions—both for perturbed cluster models and for more general modulated SPT states—remains an interesting avenue for future research. Representing a two-dimensional topological state such as the toric code and its variants [13, 19, 9], as well as the SPT states related to them [21] with ${\mathbb{Z}}_{N}$ physical degrees of freedom is another exciting avenue of future research.